# Convective Space-Time Chaos as a Dynamical Model of Deterministic and Stochastic Turbulence

### v1

@@ IV.1 Noise-sustained turbulence
We illustrate noise-sustained turbulence in the NHKS equation (1) in Fig. 4. As a noisy force, we use an Ornstein-Uhlenbeck stationary Gaussian process with zero mean and exponential autocorrelation function $C_{f}(\tau)=\langle f(t)f(t+\tau)\rangle=0.04\exp[-4\tau]$. For characterizing the statistical characteristics of the field $u(x,t)$ at different distances from the inlet, we also used an autocorrelation function $C(\tau;x)=\langle(u(x,t)-\langle u\rangle u)(u(x,t+\tau)-\langle u\rangle u)\rangle$. The spatial evolution of the random field has two features: (i) the variance grows (because the driving force has a small amplitude in this example) and saturates at large distances from the inlet; (ii) the shape of the autocorrelation function changes from monotonous exponential to oscillatory exponential. One can see from the bottom panel of Fig. 4 that the autocorrelation functions for all $x>60$ practically coincide, indicating the onset of a spatially statistically homogeneous space-time chaos.
@@ IV.2 Synchronization by common noise and reliability
As discussed in Section III, in the regimes of noise-sustained turbulence (Section IV.1), the usual temporal LE is negative. This allows for application of the concept of synchronization by common noise [31, 32, 33, 34, 35] to these regimes. This effect appears if the same noise drives two identical systems. Then, due to symmetry, a regime exists where the states of two systems coincide identically. The stability of this synchronous state is governed by the usual temporal LE. If it is negative, the synchronous state is stable, and typically establishes after some transient if the initial conditions of the two driven systems are different.
@@
$\frac{dv_{n}}{dt}=-v_{n}+F(v_{n-1}),\qquad\frac{dv_{1}}{dt}=-v_{n}+f(t)\;.$
@@
A similar simulation for the NHKS model (1) and its replica $v(x,t)$ shows that synchrony establishes only in a domain close to the inlet $x\lesssim 500$, while for larger $x$ the fields of the system $u(x,t)$ and of its replica $v(x,t)$ remain different. We will explain this feature in the next Section IV.3.
@@ IV.3 Stability of synchronous state and spatial Lyapunov exponent
Here, we illustrate how practically this exponent is calculated for the NHKS model (1). Having a space-time chaotic solution $u(x,t)$, the linearized problem is formulated as a PDE
@@
This equation is subject to initial $U(x,0)$ and to boundary conditions. We use the same boundary conditions as for the NHKS equation (cf. discussion after Eq. (3)), with some arbitrary stationary in time function $f(t)$ (e.g., one can just take the same noisy driving as used in Fig. 5(b)). Then, at different positions $x$ at large enough (so that the initial conditions move away) times, one observes statistically stationary in $t$ fields; we illustrate this in Fig. 6. Because only the norm is important, we coarse-grain the fields $U(x,t)$ by calculating
@@
One can estimate the spatial LE also from the convective LEs. Consider a perturbation $U(x,t)$ at some position $x$ and time $t$. To this perturbation contribute different perturbations initiated at the boundary $x=0$. We estimate the norm of $U(x,t)$ as the maximum over different contributions. A perturbation at the boundary at time $t-\tau$ contributes to $U(x,t)$ with the exponential factor $\exp[\lambda(v)\tau]=\exp[\lambda(v)v^{-1}x]$, where $v=x/\tau$ is the corresponding velocity. Thus, we have to maximize $\lambda(v)/v$. This condition yields the relation $\lambda^{\prime}(v^{*})v^{*}-\lambda(v^{*})=0$ for the optimal velocity. Thus, the convective exponent based estimation of the spatial exponent is
@@
Geometrically, this is the slope of the tangential to the profile of the convective LE, as shown in Fig. 2 with a dashed line. Using the expression for the convective LE in the caption of Fig. 2, we find $\Lambda\approx 0.31$, which corersponds well with the direct numerical evaluation using (14), which yields $\Lambda\approx 0.314$.
@@
Positive spatial LE explains observation of a non-perfect synchrony in the simulations of the NHKS equation Fig. 5(b). The reason is in small perturbations at the level of accuracy of the representation of double-precision numbers in numerical simulations. These perturbations are unavoidable because the coupling of different sites in the discrete representation (3) is bidirectional. Estimating the level of these perturbations as $10^{-15}$, and taking into account that for $V=2$ the spatial LE is $\approx 0.06$, we conclude that perturbations reach level one at distance $\approx\ln(10^{15})/0.06\approx 575$, what corresponds to the right border of the synchronized domain in Fig. 5(b). Exactly in the same way, positive spatial LE explains also an exponential growth of stochastic turbulence component in experiments, as is shown in Fig. (27) of Ref. [7].
We should mention, however, that the spatial LE provides, for a setup like depicted in Fig. 5(b), only an estimate of the distance at which the reliability breaks due to uncontrolled perturbations at the boundary and in the bulk. The reason is that the spatial LE is well-defined only for a spatially stationary regime like one corresponding to the bottom panel of Fig. 4. Closer to the inlet, there is a domain where statistical properties of the field are not stationary but heavily depend on the drifing force $f(t)$. Correspondingly, spatial amplification of perturbations in this non-stationary domain may deviate from that described by the spatial LE.
@@ V Periodic and quasiperiodic input
where $T_{1,2}=2\pi/\omega_{1,2}$ are two incommennsurate periods.
@@
Although the regime in the quasiperiodically driven convective system remains quasiperiodic, it becomes more and more complex downstream, and at large distances, it is practically indistinguishable from a random signal. This has been demonstrated for system (5) in [11, 12]. Below we show this spatial development of field complexity for the NHKS model (1).
@@ V.1 Evolution of a torus in the phase space
Figure 8: Images of the torus at different $x$ for $V=2$ and $T_{1}=10$. For $x=0,10,20$ only the line connecting the points is shown (because the points are extremely dense). For $x=30,40,50$ the points and the line connecting them are shown. One cannot distinguish points for $x=30$, but for $x=40,50$ one can see them because the separation between the points becomes large. For $x=60,70,80$ only points are shown; moreover, for better visibility the point size is reduced.
@@
We quantify this phenomenon of growing length of the invariant curve in Fig. 9. We present here three calculations of the length with different Fibonacci numbers of points. Reliable lengths are those where these curves coincide ($x<50$), which corresponds to the images of Fig. 8. Fig. 9 demonstrates that after some oscillations at small $x$, the length grows as $L\propto\exp[\gamma x]$ (black dashed line in Fig.9), with $\gamma\approx 0.1$. This exponent characterizes the spatial growth of the complexity of the signal and its evolution toward a random process.
@@ V.2 Evolution of the spectrum
where $w_{k}$ are normalized spectral components $w_{k}=\frac{W_{k}}{\sum_{k}W_{k}}$. The evolution of the spectral entropy for several periodic approximations $\omega_{2}/\omega_{1}=q_{m}/q_{m+1}$ where $q_{m}$ are Fibonacci numbers, is presented in Fig. 11. One can see three stages: (i) initially for $x\lesssim 5$ there is a strong growth of higher harmonics of the two initial frequencies; (ii) for $x\gtrsim 5$ there is a linear growth of entropy until a saturation, which depends on the total number of harmonics ($\mathcal{S}_{max}\propto\ln q_{m+1}$). Because the entropy is a logarithmic measure of complexity, its logarithmic growth corresponds to the exponential growth of the invariant curve length in Section V.1. Both figures 9 and 11 illustrate exponential spatial growth of complexity at a quasiperiodic driving, so that at large distances, with a finite resolution, one cannot distinguish a complex quasiperiodic process from a random one.
@@ VI Space-time duality
In the convective space-time chaos, one considers the dynamics in the domain $x\geq 0,\;t\geq 0$, subject to initial conditions at $t=0$ and boundary conditions at $x=0$. The perturbations propagate with some velocity from both boundaries into the bulk of the domain, as illustrated in Fig. 1. Due to the symmetry $t\leftrightarrow x$ of the setup, one can exchange space and time: one interprets $x$ as “time” and $t$ as “space”. Below we will keep this terminology: we will use “time” and “space” in quotation marks, if we want to denote the variable along which the evolution takes place, and the variable in which the corresponding “initial conditions” are set, respectively.
For the discrete model of a chain of amplifiers (5), one interprets $n$ as discrete time, so that the evolution in $n$ can be written as
@@
Representation (20) is mostly suitable for studying periodic in “space” $t$ fields, as such a setup is typical in studies of space-time chaos. This problem corresponds exactly to the problem of spatial evolution under periodic forcing, formulated at the beginning of Section V. The period of the force $T$ corresponds to the period in “space” in (20). The usual largest LE of (20) determines whether the evolution in “time” is chaotic or not, for large $T$ it is the same as the spatial LE (14). With this space-time duality, reliability discussed in Section IV.2 corresponds to the fact, that the “temporal” evolution from two identical “initial conditions” will be exactly the same. If, however, the “initial conditions” are slightly non-identical, then they grow exponentially and after a logarithmically short “time” the states of two runs will be different (onset of stochastic turbulence). Furthermore, one can characterize the instability related to a non-perfect periodicity of the driving with Bloch-Lyapunov exponent [11, 26], but we will not go into details.
Let us now discuss space-time duality for the NHKS equation (1). In this equation, the variables $x$ and $t$ enter highly asymmetrically, and the first derivative in $t$ indicates that evolution is along this variable. However, one can argue that for large velocities $V$, one can exchange $t$ and $x$ is an approximate way. Let us first make a change of variables $z=Vt$. Then (1) reduces to
@@
This is the original NHKS (up to transformation $u\to-u$), where the space and time are exchanged: the “time” variable $\tau$ corresponds to $x$ and the “space” variable $z$ corresponds to $t$. Such a duality is mostly simple for linear equations. Suppose one has a dispersion relation $R(\omega,k)=0$. Suppose one wants to follow the evolution of a spatial profile over time. In that case, it is natural to assume that the wavenumber $k$ is real, and to express the frequency $\omega$, which then will be generally complex, as a function (typically a polynomial) of $k$. This leads to a PDE in the form (1). If, in contradistinction, one is interested in the evolution in space of the inlet field depending on time, then it is natural to assume the frequency $\omega$ to be real, and to express the complex wavenumber $k$ as a function (typically a polynomial) of frequency $\omega$; this leads to an equation in the form (22). We note here that periodic external forcing discussed at the beginning of Section V corresponds to periodic in “space” boundary conditions in the dual representation (22).
@@
For the problem of waves on an inclined fluid flow, for which the NHKS equation (1) has been derived, the dual representation (22) makes sense for large values of parameter $V$. There is, however, a situation where a natural way to describe the dynamics is to consider the evolution along the spatial variable and treating time dependence as effective space dependence. This is the case of nonlinear optics in one-dimensional media (fibers) [41, 42], where the speed of light is certainly always large. A typical equation of the slowly varying complex amplitude $A(z,t)$ of the radiation field is a generalized nonlinear Schroedinger equation
@@ VII Discussion
    Evolution, along the spatial coordinate, of an imposed random signal toward a regime where statistical characteristics of space-time chaos do not depend on the forcing. The most straightforward interpretation of this feature is through the space-time duality (Section VI). In the dual representation, the boundary forcing corresponds to “initial conditions”, and development in space to evolution in “time”. It is known that on an initial stage, the dynamical system remembers initial conditions. However, after the trajectory converges to an attractor, initial conditions are forgotten, and one observes, in the case of a chaotic attractor, irregular chaotic dynamics, statistical properties of which do not depend on the initial conditions.
@@
Furthermore, we discussed the transformation of periodic and quasiperiodic signals in convective space-time chaos. Periodic boundary forcing remains leads to a periodic in time field downstream. If the period is large, much larger than the characteristic time scale of temporal correlations of chaos, then at any distance from the inlet, the field will have the same period, but within this period will look like a random one. This, in particular, has been observed in early experiments with the boundary layers [4], where periodic driving with a period of about 20 characteristic time scales was used. In terms of the space-time duality, this setup corresponds to periodic in “space” initial conditions, which remain periodic. A destruction of periodicity in the presence of small non-periodic perturbations is governed by the Bloch-Lyapunov exponent [26].
@@
The concept of convective space-time chaos generalizes notions of linear absolute and convective instability (in discrete situations, one often speaks about normal and non-normal matrices [43]) to secondary instabilities on top of space-time chaos. We note here that the conditions of convectivity for linear and secondary perturbations may differ. Thus, it would be interesting to explore cases where, e.g., linear instability is absolute but secondary instability is convective. One field of application of these concepts are chemical reactions in flow systems [44].

### v2

@@ IV.1 Noise-sustained turbulence
We illustrate noise-sustained turbulence in the NHKS equation (1) in Fig. 5. As a noisy force, we use an Ornstein-Uhlenbeck stationary Gaussian process with zero mean and exponential autocorrelation function $C_{f}(\tau)=\langle f(t)f(t+\tau)\rangle=0.04\exp[-4\tau]$. For characterizing the statistical characteristics of the field $u(x,t)$ at different distances from the inlet, we also used an autocorrelation function $C(\tau;x)=\langle(u(x,t)-\langle u\rangle)(u(x,t+\tau)-\langle u\rangle)\rangle$. The spatial evolution of the random field has two features: (i) the variance grows (because the driving force has a small amplitude in this example) and saturates at large distances from the inlet; (ii) the shape of the autocorrelation function changes from monotonous exponential to oscillatory exponential. One can see from the bottom panel of Fig. 5 that the autocorrelation functions for all $x>60$ practically coincide, indicating the onset of a spatially statistically homogeneous space-time chaos.
@@ IV.2 Synchronization by common noise and reliability
As discussed in Section III, in the regimes of noise-sustained turbulence (Section IV.1), the usual temporal LE is negative. This allows for application of the concept of synchronization by common noise [47, 48, 49, 50, 51] to these regimes. This effect appears if the same noise drives two identical systems. Then, due to symmetry, a regime exists where the states of two systems coincide identically. The usual temporal LE governs the stability of this synchronous state. If it is negative, the synchronous state is stable, and typically establishes after some transient if the initial conditions of the two driven systems are different.
@@
$\frac{dv_{n}}{dt}=-v_{n}+F(v_{n-1}),\qquad\frac{dv_{1}}{dt}=-v_{1}+f(t)\;.$
@@
A similar simulation for the NHKS model (1) and its replica (in terms of variable $v(x,t)$) shows that synchrony establishes only in a domain close to the inlet $x\lesssim 500$, while for larger $x$ the fields of the system $u(x,t)$ and of its replica $v(x,t)$ remain different. We will explain this feature in the next Section IV.3.
@@ IV.3 Stability of synchronous state and spatial Lyapunov exponent
Here, we illustrate how to calculate this exponent in practice, for the NHKS model (1). Having a space-time chaotic solution $u(x,t)$ (which we assume to be in a statistically stationary, typical case of space-time chaos; see however, a discussion at the end of this section), the linearized problem is formulated as a PDE
@@
This equation is subject to the initial $U(x,0)$ and to boundary conditions. We use the same boundary conditions as for the NHKS equation (cf. discussion after Eq. (3)), with some arbitrary stationary in time function $f(t)$ (e.g., one can just take the same noisy driving as used in Fig. 6(b)). Then, at different positions $x$ at large enough (so that the initial conditions move away) times, one observes statistically stationary in $t$ fields; we illustrate this in Fig. 7. Because only the norm is important, we coarse-grain the fields $U(x,t)$ by calculating
@@
One can estimate the spatial LE also from the convective LEs. Consider a perturbation $U(x,t)$ at some position $x$ and time $t$. To this perturbation contribute different components initiated at the boundary $x=0$. We estimate the norm of $U(x,t)$ as the maximum over different contributions. A perturbation at the boundary at time $t-\tau$ contributes to $U(x,t)$ with the exponential factor $\exp[\lambda(v)\tau]=\exp[\lambda(v)v^{-1}x]$, where $v=x/\tau$ is the corresponding velocity. Thus, we have to maximize $\lambda(v)/v$. This condition yields the relation $\lambda^{\prime}(v^{*})v^{*}-\lambda(v^{*})=0$ for the optimal velocity. Thus, the convective exponent-based estimation of the spatial exponent is
@@
Geometrically, this is the slope of the tangential to the profile of the convective LE, as shown for model (7) in Fig. 3 with a dashed line. For this model, using the expression for the convective LE in the caption of Fig. 3, we find $\Lambda\approx 0.31$, which corresponds well with the direct numerical evaluation using (16), which yields $\Lambda\approx 0.314$.
@@
Positive spatial LE explains observation of a non-perfect synchrony in the simulations of the NHKS equation Fig. 6(b). The reason lies in small perturbations at the level of accuracy of double-precision number representation in numerical simulations. These perturbations are unavoidable because the coupling of different sites in the discrete representation (3) is bidirectional. Estimating the level of these perturbations as $10^{-15}$, and taking into account that for $V=2$ the spatial LE is $\approx 0.06$, we conclude that perturbations reach level one at distance $\approx\ln(10^{15})/0.06\approx 575$, what corresponds to the right border of the synchronized domain in Fig. 6(b). Exactly in the same way, positive spatial LE also explains an exponential growth of stochastic turbulence component in experiments, as is shown in Fig. (27) of Ref. [30].
Figure 9: (a) Discrepancy between two identical systems (7), driven by the same turbulent signal $f(t)$; to one of the systems, a small additional Gaussian noise with standard deviations as in the markers keys is added. The dashed line has slope $\Lambda=0.314$. (b) The same data as in panel (a), but shown vs the rescaled distance $n-\frac{\ln\delta(0)}{\Lambda}$. The norm (18) is used for $\delta$.
To establish scaling properties of the growth of the stochastic component, let us denote its level along the system as $\delta(X)$, with a fixed inlet level $\delta(0)$. If $\delta(0)$ is small, growth of the stochastic component follows the exponential law (16), i.e.
$\delta(X)\sim\delta(0)\exp[\Lambda X]=\exp\left[\Lambda\left(X-\frac{\ln\delta(0)}{\Lambda}\right)\right],$ (21)
until a saturation at the level $\delta\approx 1$ occurs. Scaling expression (21) shows, that in the setup where the same input signal (up to some noisy discrepancy $\delta(0)$) is applied to two identical replicas of a convectively unstable medium, then the evolution of the difference between two systems downstream can be expressed as an function of single quantity $X-\Lambda^{-1}\ln\delta(0)$. We illustrate this in Fig. 9, where we use model (7) because it allows for a perfect control of the inlet perturbations.
We should mention, however, that the spatial LE provides, for a setup like depicted in Fig. 6(b), only an estimate of the distance at which the reliability breaks due to uncontrolled perturbations at the boundary and in the bulk. The reason is that the spatial LE is well-defined only for a spatially stationary regime like one corresponding to the bottom panel of Fig. 5. Closer to the inlet, there is a domain where statistical properties of the field are not stationary but heavily depend on the driving force $f(t)$. Correspondingly, spatial amplification of perturbations in this non-stationary domain may deviate from that described by the spatial LE.
@@ V Periodic and quasiperiodic input
where $T_{1,2}=2\pi/\omega_{1,2}$ are two incommensurate periods.
@@
Although the regime in the quasiperiodically driven convective system remains quasiperiodic, it becomes more and more complex downstream, and at large distances, it is practically indistinguishable from a random signal. This has been demonstrated for system (7) in [16, 17]. Below, we show this spatial development of field complexity for the NHKS model (1).
@@ V.1 Evolution of a torus in the phase space
Figure 10: Images of the torus at different $x$ for $V=2$ and $T_{1}=10$. For $x=0,10,20$ only the line connecting the points is shown (because the points are extremely dense). For $x=30,40,50$ the points and the line connecting them are shown. One cannot distinguish points for $x=30$, but for $x=40,50$ one can see them because the separation between the points becomes large. For $x=60,70,80$ only points are shown; moreover, for better visibility, the point size is reduced.
@@
We quantify this phenomenon of growing length of the invariant curve in Fig. 11. We present here three calculations of the length with different Fibonacci numbers of points. Reliable lengths are those where these curves coincide ($x<50$), which corresponds to the images of Fig. 10. Fig. 11 demonstrates that after some oscillations at small $x$, the length grows as $L\propto\exp[\gamma x]$ (black dashed line in Fig.11), with $\gamma\approx 0.1$. This exponent characterizes the spatial growth of the signal’s complexity and its evolution toward a random process.
@@ V.2 Evolution of the spectrum
where $w_{k}$ are normalized spectral components $w_{k}=\frac{W_{k}}{\sum_{k}W_{k}}$. The evolution of the spectral entropy for several periodic approximations $\omega_{2}/\omega_{1}=q_{m}/q_{m+1}$ where $q_{m}$ are Fibonacci numbers, is presented in Fig. 13. One can see three stages: (i) initially for $x\lesssim 5$ there is a strong growth of higher harmonics of the two initial frequencies; (ii) for $x\gtrsim 5$ there is a linear growth of entropy until a saturation, which depends on the total number of harmonics ($\mathcal{S}_{max}\propto\ln q_{m+1}$). Because the entropy is a logarithmic measure of complexity, its logarithmic growth corresponds to the exponential growth of the invariant curve length in Section V.1. Both figures 11 and 13 illustrate exponential spatial growth of complexity at a quasiperiodic driving, so that at large distances, with a finite resolution, one cannot distinguish a complex quasiperiodic process from a random one.
@@ VI Space-time duality
In the convective space-time chaos, one considers the dynamics in the domain $x\geq 0,\;t\geq 0$, subject to initial conditions at $t=0$ and boundary conditions at $x=0$. The perturbations propagate with some velocity from both boundaries into the bulk of the domain, as illustrated in Fig. 2. Due to the symmetry $t\leftrightarrow x$ of the setup, one can exchange space and time: one interprets $x$ as “time” and $t$ as “space”. Below, we will keep this terminology: we will use “time” and “space” in quotation marks, if we want to denote the variable along which the evolution takes place, and the variable in which the corresponding “initial conditions” are set, respectively.
For the discrete model of a chain of amplifiers (7), one interprets $n$ as discrete “time”, so that the evolution in $n$ can be written as
@@
Representation (24) is mostly suitable for studying periodic in “space” $t$ fields, as such a setup is typical in studies of space-time chaos. This problem corresponds exactly to the problem of spatial evolution under periodic forcing, formulated at the beginning of Section V. The period of the force $T$ corresponds to the period in “space” in (24). The usual largest LE of (24) determines whether the evolution in “time” is chaotic or not, for large $T$ it is the same as the spatial LE (16). With this space-time duality, reliability discussed in Section IV.2 corresponds to the fact, that the “temporal” evolution from two identical “initial conditions” will be exactly the same. If, however, the “initial conditions” are slightly non-identical, then the discrepancy grows exponentially and after a logarithmically short “time” the states of the two runs will be different (onset of stochastic turbulence). Furthermore, one can characterize the instability related to a non-perfect periodicity of the driving with the Bloch-Lyapunov exponent [16, 42], but we will not go into details.
Let us now discuss space-time duality for the NHKS equation (1). In this equation, the variables $x$ and $t$ enter highly asymmetrically, and the first derivative with respect to $t$ indicates that evolution is along this variable. However, one can argue that for large velocities $V$, one can exchange $t$ and $x$ in an approximate way. Let us first make a change of variables $z=Vt$. Then (1) reduces to
@@
This is the original NHKS (up to transformation $u\to-u$), where the space and time are exchanged: the “time” variable $\tau$ corresponds to $x$ and the “space” variable $z$ corresponds to $t$. Such a duality is mostly simple for linear equations. Assume that one has a dispersion relation $R(\omega,k)=0$. Suppose one wants to follow the evolution of a spatial profile over time. In that case, it is natural to assume that the wavenumber $k$ is real, and to express the frequency $\omega$, which then will be generally complex, as a function (typically a polynomial) of $k$. This leads to a PDE in the form (1). If, in contradistinction, one is interested in the evolution in space of the inlet field depending on time, then it is natural to assume the frequency $\omega$ to be real, and to express the complex wavenumber $k$ as a function (typically a polynomial) of frequency $\omega$; this leads to an equation in the form (26). We note here that periodic external forcing discussed at the beginning of Section V corresponds to periodic in “space” boundary conditions in the dual representation (26). Remarkably, an equation dual to the Korteveg - de Vries equation has been recently discussed in Ref. [57], however, without consideration of the corresponding duality of initial and boundary conditions.
@@
For the problem of waves on an inclined fluid flow, for which the NHKS equation (1) has been derived, the dual representation (26) makes sense for large values of parameter $V$. There is, however, a situation where a natural way to describe the dynamics is to consider the evolution along the spatial variable and treat time dependence as effective space dependence. This is the case of nonlinear optics in one-dimensional media (fibers) [58, 59], where the speed of light is certainly always large. A typical equation of the slowly varying complex amplitude $A(z,t)$ of the radiation field is a generalized nonlinear Schrödinger equation
@@
Finally, we mention that in the context of hydrodynamics, the space-time duality is employed in the “One-Way Navier-Stokes” approach for calculations of a compressible jet [60] and of a boundary layer transition [61].
@@ VII Discussion
    Evolution, along the spatial coordinate, of an imposed random signal toward a regime where statistical characteristics of space-time chaos do not depend on the forcing. The most straightforward interpretation of this feature is through the space-time duality (Section VI). In the dual representation, the boundary forcing corresponds to “initial conditions”, and development in space to evolution in “time”. It is known that at the initial stage, the dynamical system remembers initial conditions. However, after the trajectory converges to an attractor, initial conditions are forgotten, and one observes, in the case of a chaotic attractor, irregular chaotic dynamics, statistical properties of which do not depend on the initial conditions.
@@
Furthermore, we discussed the transformation of periodic and quasiperiodic signals in convective space-time chaos. Periodic boundary forcing leads to a periodic in time field downstream. If the period is large, much larger than the characteristic time scale of temporal correlations of chaos, then at any distance from the inlet, the field will have the same period, but within this period will look like a random one. This, in particular, has been observed in early experiments with the boundary layers [27], where periodic driving with a period of about 20 characteristic time scales was used. In terms of the space-time duality, this setup corresponds to periodic in “space” initial conditions, which remain periodic. A destruction of periodicity in the presence of small non-periodic perturbations is governed by the Bloch-Lyapunov exponent [42].
@@
The concept of convective space-time chaos generalizes notions of linear absolute and convective instability (in discrete situations, one often speaks about normal and non-normal matrices [62]) to secondary instabilities on top of space-time chaos. We note here that the conditions of convectivity for linear and secondary perturbations may differ. Thus, it would be interesting to explore cases where, e.g., linear instability is absolute but secondary instability is convective. One field of application of these concepts is chemical reactions in flow systems [10]. Another interesting question for future studies is, what happens close to the transition from convective to absolute instability, e.g., if parameter $V$ in Eq, (1) is decreased toward the critical value $V\approx v_{0}=1.4$. Such a transition also occurs in a Coutte-Taylor flow between rotating cylinders with an additional throughflow, if the throughflow velocity is varied [63, 64].
While our approach is inspired by hydrodynamic experiments, the results are not restricted to this realm, but are applicable to other cases of spatial development of chaos and turbulence, e.g., in plasma physics and electronics. Our study once again stresses the well-established importance of the inlet perturbations for the evolution of convective instabilities.
In some cases, one is not interested in this evolution, and applies the inlet perturbations mimicking a fully developed regime (see review [65] for such an approach in computational fluid dynamics). In such a setup the feature 1 is not applicable, but still, the reliability property can be followed by repeating the exact form of the inflow turbulence in several numerical experiments and comparing the outputs. Our results of Sec. IV.3 imply that in numerical experiments, even small uncontrollable round-off errors will result in a loss of reliability (repeatability) in the output. Moreover, the spatial evolution of violations of reliability allows for the estimation of the spatial Lyapunov exponent.
Another often implemented option is to apply a small periodic inlet perturbation to otherwise ideal flow; in the boundary layer studies, this has been done in experiments [24, 66] and in numerical simulations [67, 68, 69]. As discussed above, formally speaking, a periodic inlet perturbation in a convectively unstable medium results in a purely periodic response. However, such a regime is typically unstable, and this instability can be described with the Bloch-Lyapunov exponent, which generalizes the usual secondary instability approach [70] to the case of irregular in space variations of a temporally periodic field. Thus, the break of periodicity (cf. recent discussion of this phenomenon in [69]) is similar to the break of reliability. The distance at which this happens depends on the Bloch-Lyapunov exponent and the level of irregular perturbations (in the case of a boundary layer, these perturbations are discussed in Ref. [71]). Our numerical simulations of the PDE model suggest that it is extremely difficult, if altogether possible, to reduce these irregular perturbations to zero. However, it is always possible to introduce them in a controlled way at a level higher than uncontrollable round-off errors, like we did in Section IV.3.